# Superconducting integrated random access quantum memory

### v1

@@ II.3 Device fidelity calculation
We employ a fidelity metric instead of efficiency one [27] to comprehensively characterize the quantum memory performance, incorporating the effects of losses within the quantum memory cell, deviations from impedance-matching conditions, and temporal mode distortions. The measured fidelity of our device reaches 57.5% during the first release period, representing a several-fold improvement over previously reported values. Analysis reveals that $\sum_{t}\,|g(t)|^{2}=0.991$, indicating that impedance-matching imperfections constitute the dominant contribution to the infidelity. The device exhibits a characteristic fidelity decay with storage time, as illustrated in Fig. 3b, with an effective decay constant $T_{decay}=11.44\ \mu s$. This corresponds to an effective quality factor of $4.3\cdot 10^{5}$, consistent with independent measurements of the resonator’s internal quality factors by fitting complex reflection coefficient $S_{11}$ [39] (see supplementary information for details).
@@ III DISCUSSION
In summary, we present a novel architecture for superconducting multi-resonator random access quantum memory, which has been experimentally verified. This design offers significant advantages over existing implementations, requiring fewer control lines while eliminating the fidelity degradation typically caused by active coupling elements due to its decoupling from internal resonators during storage stage. The demonstrated device achieves a memory cycle time of 1.51 $\mu s$ with 57.5% fidelity during the first release period and exhibits an effective fidelity decay time of 11.44 $\mu s$. The achieved storage fidelity represents a several-fold improvement over previous reports (21% in Ref. [29] and 12% in Ref. [28]). Detailed analysis identifies impedance-matching imperfections as the primary fidelity limitation, arising from non-equidistant resonator frequencies and deviations from the designed common resonator frequency. Future work will focus on addressing these specific challenges to further enhance device performance.
@@ VII Author contributions
A.R.M and I.A.R. proposed the device architecture. A.R.M., N.S.S. and V.I.P performed the main parameters calculation and designed scheme. The technology was developed by N.S.S., D.A.M, M.V.A., M.I.T., Y.A.A. and I.A.R.. A.R.M., A.I.I, E.I.M and D.E.S. configured the experimental setup and carried out the measurements. A.R.M. and N.S.S. performed numerical simulation and analysed the experimental data. A.R.M, N.S.S and I.A.R wrote the manuscript with an assistance of all authors. A.A.V. and I.A.R supervised the project.

### v2

@@ II.3 Device fidelity calculation
We employ a fidelity metric instead of efficiency one [27] to comprehensively characterize the quantum memory performance. It incorporates the effects of losses within the quantum memory cell, deviations from impedance-matching conditions, and temporal mode distortions. The measured fidelity of our device reaches 57.5% during the first release period, representing a several-fold improvement over previously reported values. Analysis reveals that $\sum_{t}\,|g(t)|^{2}=0.991$, indicating the dominant contribution to the infidelity of impedance-matching imperfections. The device exhibits a characteristic fidelity decay with storage time, as illustrated in Fig. 3b, with an effective decay constant $T_{decay}=11.44\ \mu s$. This corresponds to an effective quality factor of $4.3\cdot 10^{5}$, consistent with independent measurements of the resonator’s internal quality factors by fitting complex reflection coefficient $S_{11}$ [39] (see supplementary information for details).
@@ III DISCUSSION
In summary, we present a novel architecture for superconducting integrated on-demand quantum memory, which has been experimentally verified. This design offers significant advantages over existing implementations, requiring fewer control lines while eliminating the fidelity degradation typically caused by active coupling elements due to its decoupling from internal resonators during the storage stage. Our device demonstrates a memory cycle time of 1.51 $\mu s$ with 57.5% fidelity during the first release period and exhibits an effective fidelity decay time of 11.44 $\mu s$. The achieved storage fidelity represents a several-fold improvement over previous reports (21% in Ref. [29] and 12% in Ref. [28]). Detailed analysis identifies impedance-matching imperfections as the primary fidelity limitation, arising from non-equidistant resonator frequencies and deviations from the designed common resonator frequency. Future work will focus on addressing these specific challenges to further enhance device performance.
@@ VII Author contributions
A.R.M and I.A.R. proposed the device architecture. A.R.M., N.S.S. and V.I.P performed the main parameters calculation and designed scheme. The technology was developed by N.S.S., D.A.M, M.V.A., Y.A.A. and I.A.R.. A.R.M., A.I.I, E.I.M and D.E.S. configured the experimental setup and carried out the measurements. A.R.M. and N.S.S. performed numerical simulation and analyzed the experimental data. A.R.M, N.S.S and I.A.R wrote the manuscript with an assistance of all authors. A.V.A. and I.A.R supervised the project.